# Properties of Immersions for Systems with Multiple Limit Sets with Implications to Learning Koopman Embeddings

### v1

@@ Appendix A Proof of the Main Theorem
Note that the function $D(F(\cdot))$ is well-defined for any $\Omega\in\widehat{\mathcal{W}^{+}}\cup\widehat{\mathcal{W}^{-}}$, since by Lemmas A.36 and B.41, for any $\Omega\in\widehat{\mathcal{W}^{+}}$ and $\Gamma\in\widehat{\mathcal{W}^{-}}$, $F(\Omega)$ and $F(\Gamma)$ are respectively an $\omega$-limit set and an $\alpha$-limit set of $\dot{z}=g(z)$ on ${\mathcal{Z}}$.
@@
Since by (T2’), for any $\xi\in{\mathcal{X}}$, the corresponding trajectory is either precompact or backward precompact in ${\mathcal{X}}$, by Lemma 2.10, any $\xi\in{\mathcal{X}}$ must belong to $D(\Omega)$ for some $\Omega\in\widehat{\mathcal{W}^{+}}\cup\widehat{\mathcal{W}^{-}}$. Thus, the set ${\mathcal{X}}=\cup_{i\in I}D(\Omega_{i})$ for a countable set $I$, where $\{\Omega_{i}\}_{i\in I}=\widehat{\mathcal{W}^{+}}\cup\widehat{\mathcal{W}^{-}}$. As a result, $F({\mathcal{X}})=\cup_{i\in I}F(D(\Omega_{i}))$. By Lemma B.42, $F(D(\Omega_{i}))$ is contained by $D(F(\Omega_{i}))$. Thus,
$F({\mathcal{X}})=\cup_{i\in I}(D(F(\Omega_{i}))\cap F({\mathcal{X}})).$ (47)
By (T1’) and Lemma B.40, $D(F(\Omega_{i}))\cap F({\mathcal{X}})$ is closed in the subspace topology induced by $F({\mathcal{X}})$ and is disjoint for all $i\in I$. So $F({\mathcal{X}})$ is a disjoint union of a countable collection of closed sets. Since ${\mathcal{X}}$ is path-connected and $F$ is continuous, $F({\mathcal{X}})$ is path-connected. Thus, by a theorem of Sierpiński (1918), only one of the sets in the collection $\{D(F(\Omega_{i}))\cap F({\mathcal{X}})\}_{i\in I}$ can be nonempty. However, since points in $F(\Omega_{i})$ are limit points of $D(F(\Omega_{i}))$ and $D(F(\Omega_{i}))$ is closed (as both $\dot{z}=g(z)$ and $\dot{z}=-g(z)$ have closed basin), the set $D(F(\Omega_{i}))\cap F({\mathcal{X}})$ contains the nonempty set $F(\Omega_{i})$, for all $i\in I$. Therefore, all the $D(F(\Omega_{i}))\cap F({\mathcal{X}})$ must be the same, and thus equal to $F({\mathcal{X}})$. By (T1’) and Lemma B.40, the limit sets $F(\Omega_{i})$ must be the same for all $i\in I$ since their domains of attraction all contain $F({\mathcal{X}})$.
Denote the only element in $\{F(\Omega_{i})\}_{i\in I}$ by $\tilde{\Omega}$. Now for any $\Omega\in\mathcal{W}^{+}\backslash\widehat{\mathcal{W}^{+}}$, we pick an arbitrary $\xi\in\Omega$. By Lemma A.36, $F(\Omega)\subseteq\omega^{+}(F(\xi))$. By the construction of $\widehat{\mathcal{W}^{+}}$ and (T2’), $\xi$ is not precompact in ${\mathcal{X}}$, and thus has to be backward precompact in ${\mathcal{X}}$. Thus, by Lemma 2.10, $\omega^{-}(\xi)$ exists and is contained by $\widehat{\mathcal{W}^{-}}$. Furthermore, by Lemma B.41, $F(\omega^{-}(\xi))=\omega^{-}(F(\xi))$. By (T1’) and Lemma B.40, since $F(\xi)\in D^{+}(\omega^{+}(F(\xi)))\cap D^{-}(\omega^{-}(F(\xi)))$, we have
$\omega^{-}(F(\xi)=\omega^{+}(F(\xi))\supseteq F(\Omega).$
Since $\omega^{-}(\xi)$ is contained in $\widehat{\mathcal{W}^{-}}$, $F(\omega^{-}(\xi))=\tilde{\Omega}$. Therefore, we have $\tilde{\Omega}\supseteq F(\Omega).$ Similarly, one can show that for any $\Gamma\in\mathcal{W}^{-}\backslash\widehat{\mathcal{W}^{-}}$, $\tilde{\Omega}\supseteq F(\Gamma)$. Thus, $\tilde{\Omega}$ is the unique maximal element in the set $\{F(\Omega)\}_{\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}}$. When (T4’) is satisfied, $\widehat{\mathcal{W}^{+}}\cup\widehat{\mathcal{W}^{-}}=\mathcal{W}^{+}\cup\mathcal{W}^{-}$. Thus, $\tilde{\Omega}$ is the only element in the set $\{F(\Omega)\}_{\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}}$. $\Box$

### v5

@@ Appendix A Proof of the Main Theorem
Now, we show that for any $\Omega,\ \Omega^{\prime}\in\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$ such that $F(\Omega)\not=F(\Omega^{\prime})$, $D(F(\Omega))$ and $D(F(\Omega^{\prime}))$ must be disjoint.
Assume that $F(\Omega)\not=F(\Omega^{\prime})$. Immediately, we have
$\begin{split}D^{+}(F(\Omega))\cap D^{+}(F(\Omega^{\prime}))=\emptyset,\\ D^{-}(F(\Omega))\cap D^{-}(F(\Omega^{\prime}))=\emptyset,\\ D^{+}(F(\Omega))\cap D^{-}(F(\Omega^{\prime}))=\emptyset,\\ D^{-}(F(\Omega))\cap D^{+}(F(\Omega^{\prime}))=\emptyset,\end{split}$ (38)
where the first two equations hold by the definition of domain of attraction, and the last two equations hold due to (T1’) and Lemma B.15. By (37) and (38), if $F(\Omega)\not=F(\Omega^{\prime})$, $D(F(\Omega))$ and $D(F(\Omega^{\prime}))$ must be disjoint.
Now we are ready to show that $F(\Omega)=F(\Omega^{\prime})$ for all $\Omega,\ \Omega^{\prime}\in\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$ under conditions (T1’)-(T3’).
@@
Since by (T2’), for any $\xi\in\mathcal{X}$, the corresponding trajectory is either precompact or backward precompact in $\mathcal{X}$, by Lemma 1, any $\xi\in\mathcal{X}$ must belong to $D(\Omega)$ for some $\Omega\in\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$. Thus, the set $\mathcal{X}=\cup_{i\in I}D(\Omega_{i})$ for a countable set $I$, where $\{\Omega_{i}\}_{i\in I}=\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$. As a result, $F(\mathcal{X})=\cup_{i\in I}F(D(\Omega_{i}))$. By Lemma B.17, $F(D(\Omega_{i}))$ is contained by $D(F(\Omega_{i}))$. Thus,
$F(\mathcal{X})=\bigcup_{i\in I}(D(F(\Omega_{i}))\cap F(\mathcal{X})).$ (39)
By (T1’), $D(F(\Omega_{i}))\cap F(\mathcal{X})$ is closed in the subspace topology induced by $F(\mathcal{X})$ for all $i\in I$. Moreover, the set $D(F(\Omega_{i}))\cap F(\mathcal{X})$ is nonempty for all $i\in I$ since points in $F(\Omega_{i})$ are limit points of $D(F(\Omega_{i}))$ and thus contained in the closed set $D(F(\Omega_{i}))$ (as both $\dot{z}=g(z)$ and $\dot{z}=-g(z)$ have closed basins).
Now suppose that $F(\Omega_{i})$ and $F(\Omega_{j})$ are not equal for some $i$ and $j$ in $I$. Then, $D(F(\Omega_{i}))$ and $D(F(\Omega_{j}))$ are nonempty and disjoint, and thus (39) implies that $F(\mathcal{X})$ is a disjoint union of a countable collection of closed sets. Since $\mathcal{X}$ is path-connected and $F$ is continuous, $F(\mathcal{X})$ is path-connected. Thus, by the main theorem of [26], only one of the sets in the collection $\{D(F(\Omega_{k}))\cap F(\mathcal{X})\}_{k\in I}$ can be nonempty. This contradicts our earlier proof that $D(F(\Omega_{i}))\cap F(\mathcal{X})$ and $D(F(\Omega_{j}))\cap F(\mathcal{X})$ are disjoint and nonempty. Thus, by contradiction, the sets $F(\Omega_{k})$ must be the same for all $k\in I$. That is, $F(\Omega)$ is the same set for all $\Omega\in\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$. $\Box$
Proof of Theorem 5.9: Recall under (T1’)-(T3’), we want to show that there exists $\Omega_{\max}\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$ such that $F(\Omega_{\max})\supseteq F(\Omega)$ for any $\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$. Note that (T2’) and Lemmas A.11 and B.16 imply that $\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$ is nonempty. Let $\Omega_{\max}$ be any element in the nonempty set $\widehat{W}^{+}\cup\widehat{W}^{-}$. For any $\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$, we show $F(\Omega)\subseteq F(\Omega_{\max})$ by considering three cases:
Case 1: $\Omega\in\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-}$. By Lemma B.18, $F(\Omega)=F(\Omega_{\max})$.
Case 2: $\Omega\in\mathcal{W}^{+}\backslash(\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-})$. Pick any $\xi\in D^{+}(\Omega)$. Since $\Omega\not\in\widehat{\mathcal{W}}^{+}$, the solution of $\xi$ cannot be precompact in $\mathcal{X}$, and thus must be backward precompact by (T2’). Thus, by Lemma 1, $\omega^{-}(\xi)$ exists and belongs to $\widehat{W}^{-}$. By Lemmas B.16 and B.18, $F(\omega^{-}(\xi))=\omega^{-}(F(\xi))=F(\Omega_{\max})$. By Lemma A.11, $\omega^{+}(F(\xi))$ is nonempty and contains $F(\Omega)$. Finally, by Lemma B.15 and (T1’), we have $\omega^{+}(F(\xi))=\omega^{-}(F(\xi))$. Therefore, we have
$F(\Omega)\subseteq\omega^{+}(F(\xi))=\omega^{-}(F(\xi))=F(\Omega_{\max}).$ (40)
Case 3: $\Omega\in\mathcal{W}^{-}\backslash(\widehat{\mathcal{W}}^{+}\cup\widehat{\mathcal{W}}^{-})$. By arguments similar to Case 2, it can be shown that $F(\Omega)\subseteq F(\Omega_{\max})$.
So far, we have shown that under conditions (T1’)-(T3’), $F(\Omega)\subseteq F(\Omega_{\max})$ for any $\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$. Finally, note that (T4’) implies $\mathcal{W}^{+}=\widehat{\mathcal{W}}^{+}$ and $\mathcal{W}^{-}=\widehat{\mathcal{W}}^{-}$, and thus for any $\Omega\in\mathcal{W}^{+}\cup\mathcal{W}^{-}$, $F(\Omega)=F(\Omega_{\max})$ by Lemma B.18. $\Box$